### v1

@@ I.2 Crystalline correspondence
Without the pairing, the full spacetime and internal symmetry of Weyl fermion is the Lorentz group with the chiral U(1) symmetry
@@
The example of a Weyl fermion deformed by a superconducting vortex can also be viewed as the boundary of a (4+1)d $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ FSPT state. To be more explicit, let us parameterize the four Euclidean coordinates of the (4+1)d FSPT bulk as $(x,y,z,w)$ and the rotational plane along the $x$-$y$ plane is parameterized by $(x,y)$. In four spatial dimensions, the rotational center (that is invariant and fixed under the $x$-$y$ plane rotation) is also a 2D plane as the $z$-$w$ plane. A “fixed-point” state for (4+1)d $C_{N}\times\mathbb{Z}_{2}$ FSPT phase is to simply put a (2+1)d $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductor on the rotation center $z$-$w$ plane. This is an example of the so-called “block-state” construction Song et al. (2017). On the (3+1)d boundary, the chiral Majorana-Weyl fermion on the rotation axis (i.e. the vortex line) is precisely the edge mode of the $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductor.
More generally, one can start from $\nu$ copies of Weyl fermions and follow the same procedure to arrive at $\nu$ chiral Majorana-Weyl fermions at the rotation axis. Below, we say that this system has $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ anomaly indexed by $\nu$. Alternatively, one may view such a system as the boundary of a (4+1)d $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ SPT state, which takes the form of a stack of $\nu$ copies of $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductors at the rotation center. In Appendix A, we derive the full classification of $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ FSPTs in (4+1)d using the block-state construction.
@@ II Construction of symmetric gapped boundary
The (3+1)d deformed theory has $\nu$ (1+1)d Majorana-Weyl fermions on the rotation axis (i.e. the $z$ axis) and each of them carries a chiral central charge $c_{-}=1/2$. In order to create a fully gapped (3+1)d boundary, we follow the strategy in Yang and Cheng (2024) and use the construction shown in Fig.1. Here the 3D $x$-$y$-$z$ boundary space is parametrized by $(x,y,z)$, taken to be the $w=0$ slice of the 4D $x$-$y$-$z$-$w$ space of (4+1)d bulk. Choose $N$ half planes all terminating at the $z$ axis, the positions of which are related to each other by $C_{N}$ rotation. For example, one of the half-planes could be the plane $y=0,x\geq 0$, all the others are obtained by $C_{N}$ rotations.
On each (2+1)d plane we decorate a (2+1)d (chiral) topological order $\cal{B}$, again all placed in a $C_{N}$-symmetric way. At the 1D rotational center $z$-axis, we have $N$ of the (1+1)d edge modes from the (2+1)d topological phases $\mathcal{B}$’s on the half-planes and the (1+1)d edge conformal field theory (CFT) from the 2D rotation center (namely the $z$-$w$ plane) in the (4+1)d bulk. We require that these edge modes together can be gapped out while preserving the $C_{N}$ symmetry.
@@
The block construction, naively, does not lead to (3+1)d TQFT boundary states since the insertion of (2+1)d (chiral) topological orders breaks the Lorentz symmetry. However, if ${\cal B}$ is a (2+1)d topological gauge theory, then one can “deform” the block construction to a TQFT by converting ${\cal B}$’s into invertible topological defects in a (3+1)d TQFT. We fill the boundary with a (3+1)d topological gauge theory and then condense the bound state of 2D and 3D gauge charges such that there are no deconfined anyons on the (2+1)d planes.
For simplicity, it suffices to assume that the gauge group is an Abelian group $G_{A}$ for all our examples. Then ${\cal B}$ has a set of bosonic or fermionic anyons carrying gauge charges of $G_{A}$. They are labeled by characters (i.e. one-dimensional representations) of $G_{A}$, the collection of which forms the group $\hat{G}_{A}\equiv\rm{Hom}(G_{A},{\rm U}(1))$, the Pontryagin dual of $G_{A}$. Importantly, these anyons braid trivially with each other, so they form a Lagrangian subgroup. Here being a Lagrangian subgroup means that every anyon NOT in $\hat{G}_{A}$ must braid nontrivially with at least one anyon in $\hat{G}_{A}$. We denote anyons from (2+1)d topological gauge theory by $b_{2d}$ for $b\in\hat{G}_{A}$. In addition, they may transform nontrivially under $C_{N}$, e.g. $C_{N}$ may act projectively on the anyons.
Then on the (3+1)d boundary, we stack a (3+1)d topological gauge theory with the same gauge group $G_{A}$ and the same group of anyons $\hat{G}_{A}$. Importantly, we require that as Abelian anyon theories, the (3+1)d theory $\hat{G_{A}}$ must be identical to the (2+1)d theory ${\cal B}$, meaning that we choose 3D gauge charges $b_{3d}$ that share the same transformation properties as $b_{2d}$ under $C_{N}$. On each $\mathcal{B}$ plane, we condense the bound state $b_{2d}\bar{b}_{3d}$ for all $b\in\hat{G_{A}}$, which effectively identifies $b_{2d}$ with $b_{3d}$. By construction, the condensation preserves the $C_{N}$ symmetry and it is expected that there is no spontaneous symmetry breaking. After the condensation, all the (2+1)d planes become codimension-1 invertible topological defects in the (3+1)d topological gauge theory, implementing the 0-form $C_{N}$ symmetry. Below we denote the invertible topological defect by $\mathsf{D}$.
To see why $\mathsf{D}$ is invertible, it is useful to consider an alternative description of the construction, adopting the “symmetry-extension” method in [Wang et al., 2018]. First, we symmetrically gap out the gapless modes by introducing invertible phases $\cal{F}$ in the block-state construction. To do this, the global symmetry $C_{N}$ may need to be extended by the “gauge group” $G_{A}$ in nontrivial ways, and consequently $\cal{F}$ is enriched by the extended symmetry group. Then to restore the physical symmetry $C_{N}$, one gauges the symmetry group $G_{A}$ in the entire (3+1)d boundary space to produce a $G_{A}$ gauge theory. The invertible state $\cal{F}$ becomes an invertible defect $\mathsf{D}$ in the gauged theory.  [^3] Before gauging, $\cal{F}$ being invertible means that there exists another state ${\cal{F}}^{-1}$, such that the stacking of $\cal{F}$ and ${\cal F}^{-1}$ can be smoothly deformed to a completely trivial state without breaking the protecting symmetry. Importantly, this deformation can be generated by a local symmetry-preserving Hamiltonian. This implies that after gauging, the defects corresponding to $\cal{F}$ and ${\cal{F}}^{-1}$ are also inverse of each other. If instead we only gauge the symmetry $G_{A}$ on the (2+1)d states, they turn into topological phases described by $\cal{B}$ in the previous construction. As we will see in Section II.3, the computations are easier to handle in this “ungauged” description.
@@
In the following gapped boundary constructions, we assume that $N$ is a power of 2 ($N=2^{p}$) for the reasons discussed in the introduction. Without loss of generality, assume that the anomaly index is $\nu\geq 0$, which means there are $\nu$ chiral Majorana-Weyl fermions on the rotation axis (namely, the $z$ axis on the (3+1)d boundary).
@@ II.1 General considerations
The topological order ${\cal B}$ must have $c_{-}=-\frac{\nu}{2N}$, so it must be non-Abelian unless $\nu$ is a multiple of $N$. Since our goal is to create a fully gapped state, it must be that $N$ copies of the topological order can have a gapped boundary to an invertible state, even before adding symmetry considerations. Mathematically, it means that ${\cal B}$ has order $N$ in the “fermionic Witt group” Davydov et al. (2011, 2010). Witt group is the equivalence classes of topological orders, with the following equivalence relation: ${\cal B}_{1}$ and ${\cal B}_{2}$ are equivalent, if and only if ${\cal B}_{1}$ and ${\cal B}_{2}$ can have a gapped interface up to the stacking of invertible states. The structure of the Witt group of fermionic topological orders is highly constrained. In particular, any non-Abelian class in the fermionic Witt group has order either 2 or infinity Davydov et al. (2010). On the other hand, from the chiral central charge, one can easily show that the order of ${\cal B}$ is a multiple of $\frac{N}{(N,\nu)}$, where $(N,\nu)$ is the greatest common divisor of $N$ and $\nu$. We thus require $\frac{N}{(N,\nu)}=2$, or $\nu=\frac{N}{2}$. If $\nu\neq 0$ or $N/2$ (mod $N$), the construction cannot work. We conjecture that in these cases, the boundary must be gapless or $C_{N}$ symmetry-breaking.
@@ II.2 $\nu=\frac{N}{2}$: non-TQFT gapped state
[…]
@@
For general even integer $N\in 2\mathbb{Z}$, one can easily adopt the $N=2$ construction. Namely, one views the system as $N/2$ identical copies of the $N=2$ boundary. That is, the $n$-th copy is rotated from the 00-th copy (namely the $N$-th copy on the $x$-$z$ plane) by a $\frac{2\pi n}{N}$ angle, spanning the $x$-$y$ plane.
@@ II.3 $\nu=N$: $\mathbb{Z}_{4}$ gauge theory
Denote by $r$ the generator of the $C_{N}$ symmetry The field content of the (1+1)d theory is listed in the following table:
@@
[…]
@@ II.3.1 $N=2^{p}$
$1\rightarrow\mathbb{Z}_{4}\rightarrow C_{4N}\rightarrow C_{N}\rightarrow 1$ (5)
@@
As illustrated in Fig.2, upon ungauging the $\mathbb{Z}_{4}$ symmetry, the (2+1)d planes are further dressed with (2+1)d Majorana FSPTs protected by $\mathbb{Z}_{4}\times\mathbb{Z}_{2}^{{\rm F}}$. Recall that the classification of $\mathbb{Z}_{4}\times\mathbb{Z}_{2}^{\rm F}$ FSPT phases in (2+1)d is $\mathbb{Z}_{8}\oplus\mathbb{Z}_{2}$  Cheng et al. (2018); Wang et al. (2017a); Guo et al. (2020).
@@
$\chi_{L,1,\alpha}\rightarrow\chi_{L,1,\alpha+1},\quad$ $1\displaystyle 1\leq\alpha\leq N-$ 1 .
@@
[…]
@@
Since $22p-$ 2 is even, one can show that $12^{2p-2}-$ 1 is a multiple of $3$, using
$x^{2k}-1=(x+1)(x^{2k-1}-x^{2k-2}+\cdots-1).$ (12)
@@
[…]
@@ II.3.2 $N=0$ mod 8
$1\rightarrow\mathbb{Z}_{2}\rightarrow C_{2N}\rightarrow C_{N}\rightarrow 1$
Denote the generator of $\mathbb{Z}_{2}$ symmetry as $g$, the group extension suggests $r^{N}=g$. As before, $\mathbb{Z}_{2}$ would eventually become the gauge symmetry. The (2+1)d planes around the rotational axis are further decorated by $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{F}$ Majorana FSPT. The generating phase has the edges modes of one left-moving Majorana-Weyl fermion $\chi_{L}$ and one right-moving Majorana-Weyl fermion $\chi_{R}$. Under the $g$ symmetry, the edge modes transform as
@@
$\begin{split}&\chi_{\alpha,L}\rightarrow\chi_{\alpha+1,L}\quad\alpha=1,\cdots,N-1\\ &\chi_{N,L}\rightarrow-\chi_{N,L}\\ &\chi_{\alpha,R}\rightarrow\chi_{\alpha+1,R}\,.\end{split}$
@@
$\psi_{l,L}\sim\sum_{\alpha=1}^{N}e^{-\frac{\pi\hskip 1.0pt\mathrm{i}\hskip 1.0pt}{N}l\alpha}\,\chi_{\alpha,L}$ $\quad l=1,3,\cdots,N-1$ (13)
@@
Suppose we have $n$ copies of the fundamental $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{F}$ Majorana FSPTs, the invariant for the $\mathbb{Z}_{2N}$ anomaly is

### v2

@@ I.2 $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ symmetry
More formally, without the pairing the full spacetime and internal symmetry of Weyl fermion is the Lorentz group with the chiral U(1) symmetry
@@
The example of a Weyl fermion deformed by a superconducting vortex can also be viewed as the boundary of a (4+1)d $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ FSPT state. To be more explicit, let us parameterize the four Euclidean coordinates of the (4+1)d FSPT bulk as $(x,y,z,w)$ and the rotational plane along the $x$-$y$ plane is parameterized by $(x,y)$. In four spatial dimensions, the rotational center (that is invariant and fixed under the $x$-$y$ plane rotation) is also a 2D plane as the $z$-$w$ plane. A “fixed-point” state for (4+1)d $C_{N}\times\mathbb{Z}_{2}$ FSPT phase is to simply put a (2+1)d $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductor on the rotation center $z$-$w$ plane. This is an example of the so-called “block-state” construction Song et al. (2017). On the (3+1)d boundary, the (1+1)d Majorana-Weyl fermion on the (1+1)d rotation axis (i.e. the vortex line) is precisely the (1+1)d edge mode of the (2+1)d $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductor.
More generally, one can start from $\nu$ copies of Weyl fermions and follow the same procedure to arrive at $\nu$ Majorana-Weyl fermions at the rotation axis. Below, we say that this system has $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ anomaly indexed by $\nu$. Alternatively, one may view such a system as the boundary of a (4+1)d $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ SPT state, which takes the form of a stack of $\nu$ copies of $p_{x}+\hskip 1.0pt\mathrm{i}\hskip 1.0ptp_{y}$ superconductors at the (2+1)d rotation center. In Appendix A, we derive the full classification of $C_{N}\times\mathbb{Z}_{2}^{\rm F}$ FSPTs in (4+1)d using the block-state construction.
@@ II Construction of symmetric gapped boundary
The (3+1)d deformed theory has $\nu$ copies of (1+1)d Majorana-Weyl fermions on the rotation axis (i.e. the $z$ axis) and each of them carries a chiral central charge $c_{-}=1/2$. In order to create a fully gapped (3+1)d boundary, we follow the strategy in Yang and Cheng (2024) and use the construction shown in Fig.1. Here the 3D $x$-$y$-$z$ boundary space is parametrized by $(x,y,z)$, taken to be the $w=0$ slice of the 4D $x$-$y$-$z$-$w$ space of (4+1)d bulk. Choose $N$ half planes all terminating at the $z$ axis, the positions of which are related to each other by $C_{N}$ rotation. For example, one of the half-planes could be the plane $y=0,x\geq 0$, all the others are obtained by $C_{N}$ rotations.
[…]
@@
[…]
@@
In the following gapped boundary constructions, we only need to focus on that $N$ is a power of 2 (namely $N=2^{p}$), for the reasons discussed in the introduction; other cases are related to the bosonic cases already explored in Yang and Cheng (2024). Without loss of generality, assume that the anomaly index is $\nu\geq 0$, which means there are $\nu$ Majorana-Weyl fermions on the rotation axis (namely, the $z$ axis on the (3+1)d boundary).
@@ II.1 General considerations
[…]
@@ II.2 $\nu=\frac{N}{2}$: non-TQFT gapped state
Let us start from $N=2$ and $\nu=\frac{N}{2}=1$. We fix an arbitrary plane passing through the rotation axis. The $C_{2}$ rotation reduces to the reflection on this plane. Exactly the same problem was studied for the surface of the crystalline topological superconductor Fidkowski et al. (2013); Cheng (2018), and one can construct a (2+1)d $\mathrm{SO}(3)_{3}\boxtimes\,\text{semion-fermion}$ topological order (TO) preserving all symmetries Cheng (2018). Recall that surface anomalies of (3+1)d crystalline topological superconductors are classified by a $\mathbb{Z}_{16}$ index $k$. The $\mathrm{SO}(3)_{3}$ Chern-Simons theory [^5] Note that the $\mathrm{SO}(3)_{3}$ Chern-Simons theory can be constructed out of the anyon contents of $\mathrm{SU}(2)_{6}$ Chern-Simons theory but only keeping its integer-spin representation ($j=0,1,2,3$) of SU(2) while removing the half-integer-spin representation ($j=1/2,3/2,5/2$) of SU(2). It should not be confused with the Fibonacci anyon theory, which can be described as the subcategory of integer-spin representations ($j=0,1$) of $\mathrm{SU}(2)_{3}$ Chern-Simons theory. In our convention, Fibonacci anyon theory is roughly the $\mathrm{SO}(3)_{\frac{3}{2}}$” Chern-Simons theory. fermionic TO carries the anomaly index $k=3$ Wang and Levin (2017); Cheng (2018); Mao and Wang (2020), and the semion-fermion TO theory carries the anomaly index 2-2 (by choosing the appropriate symmetry fractionalization on the semion) Fidkowski et al. (2013). Thus together they have anomaly index k=3+(2)=1k=3+(-2)=1. By the dimensional reduction to the plane, it is clear that $k$ and $\nu$ are in fact equal.
@@
For general even integer $N\in 2\mathbb{Z}$, one can easily adopt the $N=2$ construction. Namely, one views the system as $N/2$ identical copies of the $N=2$ boundary. That is, the $n$-th copy is rotated from the 0-th copy (namely the $N$-th copy on the $x$-$z$ plane) by a $\frac{2\pi n}{N}$ angle, spanning the $x$-$y$ plane.
@@ II.3 $\nu=N$: $\mathbb{Z}_{4}$ gauge theory
Denote the generator of the $C_{N}$ symmetry by $r$. The field content of the (1+1)d theory is listed in the following table:
@@
In order to gap out these modes, we generalize the “symmetry extension” method  Witten (2016); Wang et al. (2018) to our case. The idea is to enlarge the symmetry group so that the SPT order becomes trivialized, but then gauge the extra symmetries to recover the physical symmetry group. In this way, one obtains a topological gauge theory as the boundary state. To determine the candidate gauge group, we give the following intuition: The smallest gauge group is $\mathbb{Z}_{2}$. Suppose we extend $C_{N}$ by $\mathbb{Z}_{2}$ (i.e. ungauging the $\mathbb{Z}_{2}$ symmetry), the (2+1)d planes carry $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{\rm F}$ SPTs. Knowing that (2+1)d $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{\rm F}$ SPTs are $\Omega_{3}^{\rm Spin}(B\mathbb{Z}_{2})\cong\mathbb{Z}_{8}$ classified and any of the so-called Majorana SPTs here is of order 8. The $\mathbb{Z}_{2}$ gauge group cannot work unless $N$ is a multiple of 8. So for general $N=2^{p}$, the next candidate gauge group is $\mathbb{Z}_{4}$ (while $\mathbb{Z}_{3}$ is ruled out for obvious reasons). In the following calculations, we first check if a $\mathbb{Z}_{4}$ gauge theory is enough to saturate the anomaly for $N=2^{p}$, then for 8 |N|\,N, we check whether a $\mathbb{Z}_{2}$ gauge theory could be the minimal TQFT to saturate the anomaly (minimal in the sense of being the smallest gauge group).
@@ II.3.1 $N=2^{p}$
$1\rightarrow\mathbb{Z}_{4}\rightarrow C_{4N}\rightarrow C_{N}\rightarrow 1.$
@@
As illustrated in Fig.2, upon ungauging the $\mathbb{Z}_{4}$ symmetry, the (2+1)d planes are further dressed with (2+1)d Majorana FSPTs protected by $\mathbb{Z}_{4}\times\mathbb{Z}_{2}^{{\rm F}}$. Recall that the classification of $\mathbb{Z}_{4}\times\mathbb{Z}_{2}^{\rm F}$ FSPT phases in (2+1)d is $\Omega_{3}^{\rm Spin}(B\mathbb{Z}_{4})\cong\mathbb{Z}_{8}\oplus\mathbb{Z}_{2}$  Cheng et al. (2018); Wang et al. (2017a); Guo et al. (2020).
@@
$\chi_{L,1,\alpha}\rightarrow\chi_{L,1,\alpha+1},\quad$ $1\leq\alpha\leq N-1.$
@@
$\psi_{L,1,l}$ $\sim\sum_{\alpha=1}^{N}e^{-\frac{\pi\hskip 1.0pt\mathrm{i}\hskip 1.0pt}{N}l\alpha}\,\chi_{L,1,\alpha},\quad$ $l=1,3,\cdots,N-1.$
$\psi_{L,j,n}$ $\sim\sum_{\alpha=1}^{N}e^{-\frac{2\pi\hskip 1.0pt\mathrm{i}\hskip 1.0pt}{N}n\alpha}\,\chi_{L,j,\alpha},\quad$ $n=0,1,\cdots,N-1,\;j>1.$
${\psi}_{R,j,n}$ $\sim\sum_{\alpha=1}^{N}e^{-\frac{2\pi\hskip 1.0pt\mathrm{i}\hskip 1.0pt}{N}n\alpha}\,{\psi}_{R,j,\alpha},\quad$ $n=0,1,\cdots,N-1.$
@@
Since 2p22p-2 is even, one can show that $2^{2p-2}-1$ is a multiple of $3$, using
$x^{2k}-1=(x+1)(x^{2k-1}-x^{2k-2}+\cdots-1).$
@@
Therefore, the $\mathbb{Z}_{4N}$ symmetry is non-anomalous and all the edge modes can be fully gapped out without breaking the $\mathbb{Z}_{4N}$ symmetry. After this we gauge the $\mathbb{Z}_{4}$ symmetry to obtain a (3+1)d TQFT. Note that if we only gauge the symmetry in the (2+1)d Majorana FSPT state, the unit $\mathbb{Z}_{4}$ gauge flux becomes a non-Abelian anyon with quantum dimension $\sqrt{2}$ [^7] In fact, the topological order of the gauged Majorana FSPT can be identified as $\text{T-Pfaffian}\boxtimes\text{semion-fermion}$. In conclusion, we have constructed a (3+1)d $\mathbb{Z}_{4}$ gauge theory that saturates the $\nu=N$ anomaly.
@@ II.3.2 $N=0$ mod 8
$1\rightarrow\mathbb{Z}_{2}\rightarrow C_{2N}\rightarrow C_{N}\rightarrow 1.$
Denote the generator of $\mathbb{Z}_{2}$ symmetry as $g$, the group extension suggests $r^{N}=g$. As before, $\mathbb{Z}_{2}$ would eventually become the gauge symmetry. The (2+1)d planes around the rotational axis are further decorated by $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{\rm F}$ Majorana FSPT. The generating phase has the edges modes of one left-moving Majorana-Weyl fermion $\chi_{L}$ and one right-moving Majorana-Weyl fermion $\chi_{R}$. Under the $g$ symmetry, the edge modes transform as
@@
$\begin{split}&\chi_{\alpha,L}\rightarrow\chi_{\alpha+1,L}\quad\alpha=1,\cdots,N-1\,,\\ &\chi_{N,L}\rightarrow-\chi_{N,L}\,,\\ &\chi_{\alpha,R}\rightarrow\chi_{\alpha+1,R}\,.\end{split}$
@@
$\psi_{l,L}\sim\sum_{\alpha=1}^{N}e^{-\frac{\pi\hskip 1.0pt\mathrm{i}\hskip 1.0pt}{N}l\alpha}\,\chi_{\alpha,L}$ $\quad l=1,3,\cdots,N-1\,.$ (12)
@@
Suppose we have $n$ copies of the fundamental $\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{\rm F}$ Majorana FSPTs, the invariant for the $\mathbb{Z}_{2N}$ anomaly is